# Markovian Scale Prediction: A New Era of Visual Autoregressive Generation

### v2

@@ 1 Introduction
As a general modeling paradigm, autoregressive modeling has dominated text generation  [wan2023efficient, wan2024d2o, touvron2023llama, chung2024scaling, achiam2023gpt, lu2025easytext, shi2024fonts, shi2025wordcon] and multimodal understanding [xie2025collaboration, Chen_2024_CVPR, li2024llava]. Yet, due to suboptimal generation quality, it was not widely adopted in visual generation tasks for a long time. Recently, Visual AutoRegressive modeling (VAR) [VAR] converts next-token prediction to next-scale prediction, further unleashing the potential of the autoregressive modeling to generate high-quality visual content, making autoregression once again a mainstream modeling paradigm for visual generation [kumbong2025hmar, shao2025continuous, ren2024mvardecoupledscalewiseautoregressive, han2025infinity, hui2025autoregressive, zhang2025easycontrol]. However, VAR’s full-context dependency, requiring attention over all previous scales for next-scale prediction, can introduce certain challenges:
@@
These challenges motivate us to explore whether it is feasible to develop a visual autoregressive generation model without full-context dependency that outperforms traditional VAR in both performance and efficiency. Inspired by the concept of sufficient statistics [tishby2000informationbottleneckmethod, 7133169, alemi2019deepvariationalinformationbottleneck] in information theory, we suppose that each node in continuous chain-based propagation inherently maintains historical information to a certain extent. Appropriately distilling this into a representative dynamic state can achieve effective prediction without relying on all original historical information.
In light of this motivation, we develop Markov-VAR, a visual autoregressive generation model with Markov process. Markov-VAR refines next-scale prediction into Markovian scale prediction, where each scale prediction depends only on the current one rather than all previous scales. Chain-based unidirectional autoregressive modeling makes the current scale already encoding representative historical information for feature prediction [wiewiora2005learning, yang2022discrete, mao2020information]. Thus, we can treat each scale as a Markov state. However, non-full-context dependency inevitably discards substantial original historical information compared to full-context dependency. We establish a lightweight history compensation mechanism to enrich the Markov state for historical information. Specifically, we employ a sliding window that compresses certain previous scales within it into a compact history vector, which is then integrated with the Markov state to construct a representative dynamic state. By modeling these dynamic states as a Markov process, Markov-VAR achieves Markovian scale prediction.
Extensive experiments demonstrate Markov-VAR’s superior performance and efficiency. For image generation on ImageNet [deng2009imagenet], compared to VAR, Markov-VAR reduces FID from 3.61 to 3.23 (256×256) and remarkably decreases peak memory consumption of the computation state from 117.9GB to 19.1GB (1024×1024). As a foundation model, its performance and efficiency may become more promising when combined with other enhancement or acceleration techniques. Moreover, Markov-VAR outperforms many alternative-paradigm models or improved VAR-like variants in various comparisons. Our contributions are as follows:
@@ 2 Related Work
Early visual autoregressive generation models such as PixelRNN [van2016pixel], PixelCNN [van2016conditional] and PixelCNN++ [salimans2017pixelcnn++] achieved pixel-level conditional modeling, yet this paradigm suffers from limited long dependency and extremely slow sampling. Subsequent works [ramesh2021zero, esser2021imagebart, van2017neural, zhu2024scaling, ding2021cogviewmasteringtexttoimagegeneration, rqvae, yu2022scalingautoregressivemodelscontentrich, song2025makeanything, wan2024grid] treating image synthesis as a token prediction task relied on Vector Quantization (VQ) [van2017neural] to discretize continuous feature maps and autoregressively modeling visual token sequence. However, their raster-scan prediction order conflicts with the spatial structure of the images, leading to inferior performance compared to other alternative-paradigm models [kumari2023multi, guo2024i2v, ddim, ddpm, ho2020denoising, shen2025efficient, goodfellow2014generative, chen2025transanimate]. Recently, VAR [VAR] reformulates next-token prediction as next-scale prediction, further improving the generation quality of visual content. Some VAR-like variants, such as Infinity [han2025infinity], HART [tang2024hart] and VAR-CLIP [zhang2024var] extend VAR to text-to-image generation. Other further improved VAR-like models [li2024controlvar, guo2025fastvar, rajagopalan2025restorevar, xie2024litevar, wu2025nestedautoregressivemodels, jiao2025flexvarflexiblevisualautoregressive, InfinityStar, zhuang2025vargpt] perform various tasks, such as image super-resolution [qu2025visual], 3D Object Generation [chen2025sar3d] and image editing [wang2025training].
@@
The Markov assumption [markov1906rasprostranenie] that posits that the current state depends only on its immediate predecessor, has been widely employed for research in deep learning. Diffusion-based models [croitoru2023diffusion, zhang2025enhancing, zhu2025llada, nie2025large] have achieved remarkable success in various tasks based on this principle. Chain-based Markov modeling has been introduced to alleviate efficiency challenges of long-context modeling. MARCOS [liu2025marcos] and MCoT [yang2024markov] introduce structured memory nodes that reformulate chain-of-thought reasoning into a Markovian process for more lightweight inference. In visual autoregressive modeling, although HMAR [kumbong2025hmar] pioneeringly introduces Markov dependency to improve generation performance, it comes at the cost of increasing inference steps and the token sequence length. MVAR [zhang2025mvar] innovatively introduces spatial-Markov attention, focusing on reduce computational complexity. We argue that historical information preservation, more concise modeling, and more comprehensive scaling validation are essential for further advancing Markov visual autoregressive modeling. Therefore, we revisit these challenging trade-off problems and employ chain-based Markov modeling to develop a visual autoregressive foundation model that reconciles performance with efficiency.
@@ 3.1 Preliminary: Next-Scale Prediction
VAR [VAR] reformulates autoregressive modeling from next-token prediction to next-scale residual feature prediction, which generates images in a coarse-to-fine manner based on full-context dependency. The modeling process involves $T$ multi-scale residual features $\{R_{1},R_{2},\cdots,R_{T}\}$ defined by a size set $\{S_{1}\times S_{1},S_{2}\times S_{2},\cdots,S_{T}\times S_{T}\}$. At the $t$-th scale, VAR predicts the residual feature $R_{t}\in\mathbb{R}^{S_{t}\times S_{t}}$ based on all previous residual scales $R_{<t}=\{R_{1},R_{2},\cdots,R_{t-1}\}$. The autoregressive likelihood is formulated as follow:
@@ 3.3 Markov-VAR
Classical information theory [tishby2000informationbottleneckmethod] and its applications in deep learning [7133169, alemi2019deepvariationalinformationbottleneck] indicate that: information $I(c_{<t};c_{t})$ of the complete history $c_{<t}$ with the current time $c_{t}$ is highly redundant. There exists a sufficient statistic $c_{t-1}$ such that: $I(c_{t-1};c_{t})=I(c_{<t};c_{t})$. Therefore, we argue that continuous chain-based modeling can propagate this sufficient statistic, enabling effective prediction of subsequent scales only based on the current scale. This theory allows us to naturally treat each scale as a Markov state.
@@ 3.4 Training Strategy
Algorithm 1 shows the training process, where modules with subscript $\theta$ are learnable. $\hat{f}$ denotes the final target feature of Markov-VAR. $\mathcal{F}=[\langle\mathrm{sos}\rangle,X_{1},\cdots,X_{T-1}]$ represents a queue of token sequences. The overall scale sequence $\hat{R}$ consist of the residual features predicted for each scale. Finally, the training loss $\mathcal{L}$ is calculated by the cross-entropy (CE) [shannon1948mathematical] between the ground-truth $R$ with $\hat{R}$.
@@ 4.1 Experiment Settings
We construct Markov-VAR models with depths $d$ of 16, 20, and 24 layers, the network structure is set as: the width $w=64d$, the number of attention heads $h=d$ and the dropout rate $dr=0.1\cdot d/24$. We utilize Rotary Positional Embedding [su2023roformerenhancedtransformerrotary] for learnable positional encoding, and adopt LLaMA-style [touvron2023llama] attention and MLP blocks following [tang2024hart]. For image tokenization, we employ the pre-trained multi-scale VQ-VAE tokenizer from VAR. To evaluate the effectiveness of our design, we compare Marokov-VAR with the VAR [VAR] and other VAR-like models on comparable model sizes. We further include comparisons with broad generative models of alternative paradigms, such as diffusion models, GANs, and next-token prediction autoregressive models, to provide a comprehensive evaluation.
@@
We conduct the pre-training of Markov-VAR on the ImageNet-1K dataset [deng2009imagenet], following the standard class-to-image generation model setting. For quantitative evaluation, we use Fréchet Inception Distance (FID), Inception Score (IS), Precision, and Recall as the primary metrics to comprehensively assess generation quality. We evaluate efficiency in terms of inference time and peak GPU memory consumption of computation state, which is mainly composed of Transformer activations and KV cache.
@@
All models are trained in similar settings, with a base learning rate of $8\times 10^{-5}$, using the AdamW optimizer [loshchilov2017decoupled] with $\beta_{1}=0.9$ and $\beta_{2}=0.95$. The batch size ranges from 768 to 1536, and the training epochs vary from 200 to 400, depending on the model depth. Training is implemented on 8 NVIDIA H200 GPUs. Evaluation is conducted on a single NVIDIA H200 GPU.
@@
M-VAR-d20 [ren2024mvardecoupledscalewiseautoregressive] 900M 2.41 308.4 0.85 0.58
FlexVAR-d24 [jiao2025flexvarflexiblevisualautoregressive] 1.0B 2.21 299.1 0.83 0.59
HMAR-d24 [kumbong2025hmar] 1.3B 2.10 324.3 0.83 0.60
NestAR-H [wu2025nestedautoregressivemodels] 1.3B 2.22 342.4 0.79 0.57
VAR-d16[VAR] 310M 3.61 225.6 0.81 0.52
VAR-d20[VAR] 600M 2.67 254.4 0.81 0.57
VAR-d24[VAR] 1.0B 2.17 271.9 0.81 0.59
VAR-d30[VAR] 2.0B 2.14 275.4 0.80 0.60
@@
GAN BigGAN [bigGAN] 112M 6.95 224.5 0.89 0.38 1
GAN GigaGAN [kang2023scaling] 569M 3.45 225.5 0.84 0.61 1
GAN StyleGAN-XL [sauer2022stylegan] 166M 2.30 265.1 0.78 0.53 1
Diffusion ADM [dhariwal2021diffusion] 554M 10.94 101.0 0.69 0.63 250
Diffusion CDM [ho2022cascaded] - 4.88 158.7 - - 8100
Diffusion LDM-4 [rombach2022high] 400M 3.60 247.7 - - 250
Diffusion DiT-XL/2 [peebles2023scalable] 675M 2.27 278.2 0.83 0.57 250
Masked AR MaskGIT [chang2022maskgit] 227M 6.18 182.1 0.80 0.51 8
Masked AR MaskGIT-re [li2023mage] 227M 4.02 355.6 - 8
Masked AR MAGE [li2024autoregressiveimagegenerationvector] 230M 6.93 195.8 - - 20
Next-token AR VQGAN [esser2021taming] 227M 18.65 80.4 0.78 0.26 256
Next-token AR VQGAN  [esser2021taming] 1.4B 15.76 74.3 - - 256
Next-token AR VQGAN-re [yu2021vector] 1.4B 5.20 280.3 - - 256
Next-token AR ViT-VQGAN [yu2021vector] 1.7B 4.17 175.1 - - 1024
Next-token AR ViT-VQGAN-re [yu2021vector] 1.7B 3.04 227.4 - - 1024
Next-token AR RQTran [lee2022autoregressive] 3.8B 7.55 80.4 0.78 0.26 68
Next-token AR RQTran-re [lee2022autoregressive] 3.8B 3.80 323.7 - - 68
Next-token AR LlamaGen-B [llamagen]  [^1] M 5.46 193.6 0.83 0.45 256
Next-token AR LlamaGen-L [llamagen] 343M 3.81 248.3 0.83 0.52 256
Next-token AR LlamaGen-XL [llamagen] 775M 3.39 227.1 0.81 0.54 256
Next-token AR LlamaGen-XXL [llamagen] 1.4B 3.09 253.6 0.83 0.53 256
@@ 4.2 Main Results
We first evaluate the generation performance of Markov-VAR. We conduct comparsion with VAR and other VAR-like models [han2025infinity, ren2024mvardecoupledscalewiseautoregressive, kumbong2025hmar, wu2025nestedautoregressivemodels] on ImageNet $256\times 256$ class-conditional image generation. As shown in Table 1, Markov-VAR demonstrates a comprehensively better performance, achieving higher generation quality than other VAR models of comparable size. For example, on the depth-16 model, Markov-VAR improve the FID performance from 3.61 to 3.23 (10.5%) and IS performance from 225.6 to 256.2 (13.6%) compared to VAR. Even on Precision and Recall metrics, Markov-VAR matches or surpasses VAR. Markov-VAR’s advantages can be observed across other sizes. Compared with M-VAR-d20, Markov-VAR-d20 achieves compelling performance while using only 70% of its parameters, demonstrating that Markov-VAR is both effective and parameter-efficient.
To further demonstrate the performance advancements of Markov-VAR. We conduct a board comparison on class-to-image generation on ImageNet $256\times 256$ benchmark with more alternative-paradigm models, including GANs [bigGAN, kang2023scaling, sauer2022stylegan], diffusion models [dhariwal2021diffusion, ho2020denoising, rombach2022high, peebles2023scalable], mask autoregressive models [chang2022maskgit, li2023mage, li2024autoregressiveimagegenerationvector], and next-token prediction autoregressive models [llamagen, esser2021taming, yu2021vector, lee2022autoregressive]. As shown in Table 2, Compared with Diffusion models, Markov-VAR achieves better overall performance in terms of both parameter efficiency and generation quality. Compared to the size-similar next-token autoregressive model LlamaGen-L, Markov-VAR achieves superior performance across all evaluation metrics. In contrast to other autoregressive models, Markov-VAR maintains competitive generation quality while requiring fewer sampling steps, demonstrating its higher inference efficiency. Further comparisons in the table also reveal Markov-VAR’s favorable generation performance and parameter efficiency.
@@
VAR [VAR] 16 256 0.303 5.8
VAR [VAR] 16 512 0.824 18.4
VAR [VAR] 16 1024 3.303 66.7
HMAR [kumbong2025hmar] 16 256 0.309 3.1
HMAR [kumbong2025hmar] 16 512 0.909 5.6
FlexVAR [jiao2025flexvarflexiblevisualautoregressive] 16 256 0.395 6.8
FlexVAR [jiao2025flexvarflexiblevisualautoregressive] 16 512 0.944 20.1
FlexVAR [jiao2025flexvarflexiblevisualautoregressive] 16 1024 3.409 77.9

### v3

@@ 1 Introduction
As a general modeling paradigm, autoregressive modeling has dominated text generation  [66, 67, 61, 8, 1, 36, 52, 51] and multimodal understanding [71, 7, 29]. Yet, due to suboptimal generation quality, it was not widely adopted in visual generation tasks for a long time. Recently, Visual AutoRegressive modeling (VAR) [58] converts next-token prediction to next-scale prediction, further unleashing the potential of the autoregressive modeling to generate high-quality visual content, making autoregression once again a mainstream modeling paradigm for visual generation [26, 49, 44, 18, 22, 80]. However, VAR’s full-context dependency, requiring attention over all previous scales for next-scale prediction, can introduce certain challenges:
@@
These challenges motivate us to explore whether it is feasible to develop a visual autoregressive generation model without full-context dependency that outperforms traditional VAR in both performance and efficiency. Inspired by the concept of sufficient statistics [59, 60, 2] in information theory, we suppose that each node in continuous chain-based propagation inherently maintains historical information to a certain extent. Appropriately distilling this into a representative dynamic state can achieve effective prediction without relying on all original historical information.
In light of this motivation, we develop Markov-VAR, a visual autoregressive generation model with Markov process. Markov-VAR refines next-scale prediction into Markovian scale prediction, where each scale prediction depends only on the current one rather than all previous scales. Chain-based unidirectional autoregressive modeling makes the current scale already encoding representative historical information for feature prediction [69, 73, 37]. Thus, we can treat each scale as a Markov state. However, non-full-context dependency inevitably discards substantial original historical information compared to full-context dependency. We establish a lightweight history compensation mechanism to enrich the Markov state for historical information. Specifically, we employ a sliding window that compresses certain previous scales within it into a compact history vector, which is then integrated with the Markov state to construct a representative dynamic state. By modeling these dynamic states as a Markov process, Markov-VAR achieves Markovian scale prediction.
Extensive experiments demonstrate Markov-VAR’s superior performance and efficiency. For image generation on ImageNet [10], compared to VAR, Markov-VAR reduces FID from 3.61 to 3.23 (256×256) and remarkably decreases peak memory consumption of the computation state from 117.9GB to 19.1GB (1024×1024). As a foundation model, its performance and efficiency may become more promising when combined with other enhancement or acceleration techniques. Moreover, Markov-VAR outperforms many alternative-paradigm models or improved VAR-like variants in various comparisons. Our contributions are as follows:
@@ 2 Related Work
Early visual autoregressive generation models such as PixelRNN [63], PixelCNN [62] and PixelCNN++ [46] achieved pixel-level conditional modeling, yet this paradigm suffers from limited long dependency and extremely slow sampling. Subsequent works [43, 13, 64, 82, 12, 28, 76, 54, 65] treating image synthesis as a token prediction task relied on Vector Quantization (VQ) [64] to discretize continuous feature maps and autoregressively modeling visual token sequence. However, their raster-scan prediction order conflicts with the spatial structure of the images, leading to inferior performance compared to other alternative-paradigm models [25, 17, 53, 19, 20, 50, 15, 5]. Recently, VAR [58] reformulates next-token prediction as next-scale prediction, further improving the generation quality of visual content. Some VAR-like variants, such as Infinity [18], HART [57] and VAR-CLIP [78] extend VAR to text-to-image generation. Other further improved VAR-like models [32, 16, 42, 72, 70, 23, 34, 83] perform various tasks, such as image super-resolution [41], 3D Object Generation [6] and image editing [68].
@@
The Markov assumption [38] that posits that the current state depends only on its immediate predecessor, has been widely employed for research in deep learning. Diffusion-based models [9, 79, 81, 39] have achieved remarkable success in various tasks based on this principle. Chain-based Markov modeling has been introduced to alleviate efficiency challenges of long-context modeling. MARCOS [33] and MCoT [74] introduce structured memory nodes that reformulate chain-of-thought reasoning into a Markovian process for more lightweight inference. In visual autoregressive modeling, although HMAR [26] pioneeringly introduces Markov dependency to improve generation performance, it comes at the cost of increasing inference steps and the token sequence length. MVAR [77] innovatively introduces spatial-Markov attention, focusing on reduce computational complexity. We argue that historical information preservation, more concise modeling, and more comprehensive scaling validation are essential for further advancing Markov visual autoregressive modeling. Therefore, we revisit these challenging trade-off problems and employ chain-based Markov modeling to develop a visual autoregressive foundation model that reconciles performance with efficiency.
@@ 3.1 Preliminary: Next-Scale Prediction
VAR [58] reformulates autoregressive modeling from next-token prediction to next-scale residual feature prediction, which generates images in a coarse-to-fine manner based on full-context dependency. The modeling process involves $T$ multi-scale residual features $\{R_{1},R_{2},\cdots,R_{T}\}$ defined by a size set $\{S_{1}\times S_{1},S_{2}\times S_{2},\cdots,S_{T}\times S_{T}\}$. At the $t$-th scale, VAR predicts the residual feature $R_{t}\in\mathbb{R}^{S_{t}\times S_{t}}$ based on all previous residual scales $R_{<t}=\{R_{1},R_{2},\cdots,R_{t-1}\}$. The autoregressive likelihood is formulated as follow:
@@ 3.3 Markov-VAR
Classical information theory [59] and its applications in deep learning [60, 2] indicate that: information $I(c_{<t};c_{t})$ of the complete history $c_{<t}$ with the current time $c_{t}$ is highly redundant. There exists a sufficient statistic $c_{t-1}$ such that: $I(c_{t-1};c_{t})=I(c_{<t};c_{t})$. Therefore, we argue that continuous chain-based modeling can propagate this sufficient statistic, enabling effective prediction of subsequent scales only based on the current scale. This theory allows us to naturally treat each scale as a Markov state.
@@ 3.4 Training Strategy
Algorithm 1 shows the training process, where modules with subscript $\theta$ are learnable. $\hat{f}$ denotes the final target feature of Markov-VAR. $\mathcal{F}=[\langle\mathrm{sos}\rangle,X_{1},\cdots,X_{T-1}]$ represents a queue of token sequences. The overall scale sequence $\hat{R}$ consist of the residual features predicted for each scale. Finally, the training loss $\mathcal{L}$ is calculated by the cross-entropy (CE) [48] between the ground-truth $R$ with $\hat{R}$.
@@ 4.1 Experiment Settings
We construct Markov-VAR models with depths $d$ of 16, 20, and 24 layers, the network structure is set as: the width $w=64d$, the number of attention heads $h=d$ and the dropout rate $dr=0.1\cdot d/24$. We utilize Rotary Positional Embedding [55] for learnable positional encoding, and adopt LLaMA-style [61] attention and MLP blocks following [57]. For image tokenization, we employ the pre-trained multi-scale VQ-VAE tokenizer from VAR. To evaluate the effectiveness of our design, we compare Marokov-VAR with the VAR [58] and other VAR-like models on comparable model sizes. We further include comparisons with broad generative models of alternative paradigms, such as diffusion models, GANs, and next-token prediction autoregressive models, to provide a comprehensive evaluation.
@@
We conduct the pre-training of Markov-VAR on the ImageNet-1K dataset [10], following the standard class-to-image generation model setting. For quantitative evaluation, we use Fréchet Inception Distance (FID), Inception Score (IS), Precision, and Recall as the primary metrics to comprehensively assess generation quality. We evaluate efficiency in terms of inference time and peak GPU memory consumption of computation state, which is mainly composed of Transformer activations and KV cache.
@@
All models are trained in similar settings, with a base learning rate of $8\times 10^{-5}$, using the AdamW optimizer [35] with $\beta_{1}=0.9$ and $\beta_{2}=0.95$. The batch size ranges from 768 to 1536, and the training epochs vary from 200 to 400, depending on the model depth. Training is implemented on 8 NVIDIA H200 GPUs. Evaluation is conducted on a single NVIDIA H200 GPU.
@@
M-VAR-d20 [44] 900M 2.41 308.4 0.85 0.58
FlexVAR-d24 [23] 1.0B 2.21 299.1 0.83 0.59
HMAR-d24 [26] 1.3B 2.10 324.3 0.83 0.60
NestAR-H [70] 1.3B 2.22 342.4 0.79 0.57
VAR-d16[58] 310M 3.61 225.6 0.81 0.52
VAR-d20[58] 600M 2.67 254.4 0.81 0.57
VAR-d24[58] 1.0B 2.17 271.9 0.81 0.59
VAR-d30[58] 2.0B 2.14 275.4 0.80 0.60
@@
GAN BigGAN [3] 112M 6.95 224.5 0.89 0.38 1
GAN GigaGAN [24] 569M 3.45 225.5 0.84 0.61 1
GAN StyleGAN-XL [47] 166M 2.30 265.1 0.78 0.53 1
Diffusion ADM [11] 554M 10.94 101.0 0.69 0.63 250
Diffusion CDM [21] - 4.88 158.7 - - 8100
Diffusion LDM-4 [45] 400M 3.60 247.7 - - 250
Diffusion DiT-XL/2 [40] 675M 2.27 278.2 0.83 0.57 250
Masked AR MaskGIT [4] 227M 6.18 182.1 0.80 0.51 8
Masked AR MaskGIT-re [30] 227M 4.02 355.6 - 8
Masked AR MAGE [31] 230M 6.93 195.8 - - 20
Next-token AR VQGAN [14] 227M 18.65 80.4 0.78 0.26 256
Next-token AR VQGAN  [14] 1.4B 15.76 74.3 - - 256
Next-token AR VQGAN-re [75] 1.4B 5.20 280.3 - - 256
Next-token AR ViT-VQGAN [75] 1.7B 4.17 175.1 - - 1024
Next-token AR ViT-VQGAN-re [75] 1.7B 3.04 227.4 - - 1024
Next-token AR RQTran [27] 3.8B 7.55 80.4 0.78 0.26 68
Next-token AR RQTran-re [27] 3.8B 3.80 323.7 - - 68
Next-token AR LlamaGen-B [56]  [^1] M 5.46 193.6 0.83 0.45 256
Next-token AR LlamaGen-L [56] 343M 3.81 248.3 0.83 0.52 256
Next-token AR LlamaGen-XL [56] 775M 3.39 227.1 0.81 0.54 256
Next-token AR LlamaGen-XXL [56] 1.4B 3.09 253.6 0.83 0.53 256
@@ 4.2 Main Results
We first evaluate the generation performance of Markov-VAR. We conduct comparsion with VAR and other VAR-like models [18, 44, 26, 70] on ImageNet $256\times 256$ class-conditional image generation. As shown in Table 1, Markov-VAR demonstrates a comprehensively better performance, achieving higher generation quality than other VAR models of comparable size. For example, on the depth-16 model, Markov-VAR improve the FID performance from 3.61 to 3.23 (10.5%) and IS performance from 225.6 to 256.2 (13.6%) compared to VAR. Even on Precision and Recall metrics, Markov-VAR matches or surpasses VAR. Markov-VAR’s advantages can be observed across other sizes. Compared with M-VAR-d20, Markov-VAR-d20 achieves compelling performance while using only 70% of its parameters, demonstrating that Markov-VAR is both effective and parameter-efficient.
To further demonstrate the performance advancements of Markov-VAR. We conduct a board comparison on class-to-image generation on ImageNet $256\times 256$ benchmark with more alternative-paradigm models, including GANs [3, 24, 47], diffusion models [11, 20, 45, 40], mask autoregressive models [4, 30, 31], and next-token prediction autoregressive models [56, 14, 75, 27]. As shown in Table 2, Compared with Diffusion models, Markov-VAR achieves better overall performance in terms of both parameter efficiency and generation quality. Compared to the size-similar next-token autoregressive model LlamaGen-L, Markov-VAR achieves superior performance across all evaluation metrics. In contrast to other autoregressive models, Markov-VAR maintains competitive generation quality while requiring fewer sampling steps, demonstrating its higher inference efficiency. Further comparisons in the table also reveal Markov-VAR’s favorable generation performance and parameter efficiency.
@@
VAR [58] 16 256 0.303 5.8
VAR [58] 16 512 0.824 18.4
VAR [58] 16 1024 3.303 66.7
HMAR [26] 16 256 0.309 3.1
HMAR [26] 16 512 0.909 5.6
FlexVAR [23] 16 256 0.395 6.8
FlexVAR [23] 16 512 0.944 20.1
FlexVAR [23] 16 1024 3.409 77.9